The perimeter of a rectangular room is 94ft. The length is two feet longer than twice the width. Find the length of this room.
Answer by stanbon(75887)

The perimeter of a rectangular room is 94ft. The length is two feet longer than twice the width. Find the length of this room.
----------------
Perimeter = 2(L + W)
---
Let L = 2x + 2
Then W = x
--------------
Substitute:
94 = 2(2x+2 + x)
47 = 3x + 2
3x = 45
x = 15 ft. (width of the room)
2x+2 = 32 ft. (length of the room)
======================================
Cheers,
Stan H.